OpenStudy (anonymous):

Simplify radical expression.. (6 - sqaure root of 11) (6 + square root of 11)

OpenStudy (lgbasallote):

difference of two squares :DDD \[\huge (6 - \sqrt{11})(6 + \sqrt{11}) \implies (6)^2 - (\sqrt{11})^2\]